# Puzzled by Puzzles: When Vision-Language Models Can’t Take a Hint

### v1

@@ Abstract
Rebus puzzles, visual riddles that encode language through imagery, spatial arrangement, and symbolic substitution, pose a unique challenge to current vision-language models (VLMs). Unlike traditional image captioning or question answering tasks, rebus solving requires multi-modal abstraction, symbolic reasoning, and a grasp of cultural, phonetic and linguistic puns. In this paper, we investigate the capacity of contemporary VLMs to interpret and solve rebus puzzles by constructing a hand-generated and annotated benchmark of diverse English-language rebus puzzles, ranging from simple pictographic substitutions to spatially-dependent cues ("head" over "heels"). We analyze how different VLMs perform, and our findings reveal that while VLMs exhibit some surprising capabilities in decoding simple visual clues, they struggle significantly with tasks requiring abstract reasoning, lateral thinking, and understanding visual metaphors.
@@ 1 Introduction & Background
Towards solving such issues with standard benchmarking, “puzzle-based” evaluations have been explored as an alternative test for higher-order AI reasoning. In VLMs, early studies tested models on games such as Iconary Clark et al. (2021) and IQ-style challenges Huang et al. (2023), with recent benchmarks such as VGRP-Bench Ren et al. (2025) using visual grid logic puzzles to demonstrate that multi-modal language models struggle with systematic deduction and TruthQuest Mondorf and Plank (2024) using “Knights and Knaves” logic puzzles to test LLM suppositional reasoning. In textual riddles, benchmarks like RiddleSense Lin et al. (2021) test basic reasoning, while BRAINTEASER Jiang et al. (2023) and LatEval Huang et al. (2024) focus on lateral thinking, showing models often default to incorrect commonsense answers or struggle with interactive inquiry.
@@ 2 Dataset & Metrics
To evaluate the performance of VLMs, we adopt two evaluation methods: “Naive Matching” and “LLM-Judged” evaluation. The Naive Matching method evaluates responses based on exact string matching with the ground-truth answers, while the LLM-Judged evaluation method uses gpt-4o to compare ground truth and candidate answers directly (see Appendix H for details).
@@ 3 Results & Discussion
Overall performance on the probe dataset is given in Table 1. We can see that models have a wide range of performance - while closed source reasoning models such as o3, o4-mini, and gemini-2.5-pro perform relatively well, especially compared to human non-expert/non-native English speaking solvers, open-source reasoning models as non-reasoning models struggle to solve the tasks. Compared to expert solvers, however, there remains a significant model accuracy gap.
@@
SigLIP 2/So400m-16/256 29.2 43.5 29.2 36.1 46.6
SigLIP 2/So400m-16/384 28.9 43.3 28.9 36.1 46.6
SigLIP 2/So400m-16/512 28.5 41.9 28.5 35.4 46.0
@@ Acknowledgements
We would like to express our sincere gratitude to Lisa Dunlap, XuDong Wang, Konpat Preechakul, Baifeng Shi, and Stephanie Murphy for their invaluable assistance in paper review, ideation, and puzzle solving.
As part of their affiliation with UC Berkeley, the authors were supported in part by the National Science Foundation, the Ford Foundation, and/or the Berkeley Artificial Intelligence Research (BAIR) Industrial Alliance program. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA), the Army Contracting Command-Aberdeen Proving Grounds (ACC-APG), and the Air Force Research Laboratory. The views, opinions, and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of any supporting entity, including the AFRL, ACC-APG, the Department of Defense or the U.S. Government.
@@ Appendix G Image Retrieval
SigLIP 2/So400m-16/256 29.2 43.5 29.2 36.1 46.6
SigLIP 2/So400m-16/384 28.9 43.3 28.9 36.1 46.6
SigLIP 2/So400m-16/512 28.5 41.9 28.5 35.4 46.0
@@ Appendix H Evaluation Method
Using the JSON evaluation log created, we perform repeated sampling with replacement to simulate multiple evaluation scenarios. Specifically, we draw N bootstrap samples of size |S| (equal to the number of total examples), calculate the accuracy for each sample, and estimate the distribution of accuracies. Formally, for each bootstrap sample S with size |S|, the accuracy is computed as:
@@ Appendix I Prompting
Table I.1: Impact of changing the base prompt on the overall model accuracy (success).
@@
These models are capable of multi-round chain-of-thought reasoning using special "thinking" tokens, such as OpenAI o3 and o4-mini OpenAI (2025) models. Gemini-2.5-flash and Gemini-2.5-pro Kavukcuoglu (2025), both from Google, also support this functionality. Notably, Gemini 2.5-flash has both reasoning and non-reasoning variants and we use the reasoning-enabled version in our experiments.
@@
In this section, Figure K.1 shows several qualitative examples from two models (GPT-O3, Phi-4 and Qwen 2.5-VL) on our probe dataset.

### v2

@@ Abstract
Rebus puzzles, visual riddles that encode language through imagery, spatial arrangement, and symbolic substitution, pose a unique challenge to current vision-language models (VLMs). Unlike traditional image captioning or question answering tasks, rebus solving requires multi-modal abstraction, symbolic reasoning, and a grasp of cultural, phonetic and linguistic puns. In this paper, we investigate the capacity of contemporary VLMs to interpret and solve rebus puzzles by constructing a hand-generated and annotated benchmark of diverse English-language rebus puzzles, ranging from simple pictographic substitutions to spatially-dependent cues (”head” over ”heels”). We analyze how different VLMs perform, and our findings reveal that while VLMs exhibit some surprising capabilities in decoding simple visual clues, they struggle significantly with tasks requiring abstract reasoning, lateral thinking, and understanding visual metaphors.
@@ 1 Introduction & Background
Towards solving such issues with standard benchmarking, “puzzle-based” evaluations have been explored as an alternative test for higher-order AI reasoning. In VLMs, early studies tested models on games such as Iconary Clark et al. (2021) and IQ-style challenges Huang et al. (2023), with recent benchmarks such as VGRP-Bench Ren et al. (2025) using visual grid logic puzzles to demonstrate that multi-modal language models struggle with systematic deduction and TruthQuest Mondorf and Plank (2024) using “Knights and Knaves” logic puzzles to test LLM suppositional reasoning. In textual riddles, benchmarks like RiddleSense Lin et al. (2021) test basic reasoning, while BRAINTEASER Jiang et al. (2023) and LatEval Huang et al. (2024) focus on lateral thinking, showing models often default to incorrect commonsense answers or struggle with interactive inquiry. Bongard in Wonderland Dai et al. (2024) tests puzzles that require basic human-like abilities, revealing that VLMs struggle even to detect basic visual concepts and perform abstract reasoning.
@@ 2 Dataset & Metrics
To evaluate the performance of VLMs, we adopt two evaluation methods: “Naive Matching” and “LLM-Judged” evaluation. The Naive Matching method evaluates responses based on exact string matching with the ground-truth answers, while the LLM-Judged evaluation method uses Prompted LLMs (gpt-4o, Qwen3-8B) to compare ground truth and candidate answers directly (see Appendix H for details).
@@
GPT-5 $60.42_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.556,0.651]}}$ $69.44_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.653,0.736]}}$
@@
Molmo-72B-0924 $1.86_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.007,0.032]}}$ $4.43_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.025,0.065]}}$
@@
Molmo-7B-D-0924 $1.35_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.004,0.025]}}$ $1.85_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.001,0.030]}}$
@@ 3 Results & Discussion
Overall performance on the probe dataset is given in Table 1. We can see that models have a wide range of performance - while closed source reasoning models such as GPT-5, o3, and o4-mini perform relatively well, especially compared to human non-expert/non-native English speaking solvers, open-source reasoning models as non-reasoning models struggle to solve the tasks. Compared to expert solvers, however, there remains a significant model accuracy gap.
@@
SigLIP 2/So400m‑16/256 29.2 43.5 29.2 36.1 46.6
SigLIP 2/So400m‑16/384 28.9 43.3 28.9 36.1 46.6
SigLIP 2/So400m‑16/512 28.5 41.9 28.5 35.4 46.0
@@ Acknowledgements
We would like to express our sincere gratitude to Lisa Dunlap, XuDong Wang, Konpat Preechakul, Baifeng Shi, and Stephanie Murphy for their invaluable assistance in paper review, ideation, and puzzle solving. As part of their affiliation with UC Berkeley, the authors were supported in part by the National Science Foundation, the Ford Foundation, and/or the Berkeley Artificial Intelligence Research (BAIR) Industrial Alliance program. This material is based upon work supported by the Defense Advanced Research Projects Agency (DARPA), the Army Contracting Command-Aberdeen Proving Grounds (ACC-APG), and the Air Force Research Laboratory. The views, opinions, and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of any supporting entity, including the AFRL, ACC-APG, the Department of Defense or the U.S. Government.
@@ Appendix G Image Retrieval
SigLIP 2/So400m‑16/256 29.2 43.5 29.2 36.1 46.6
SigLIP 2/So400m‑16/384 28.9 43.3 28.9 36.1 46.6
SigLIP 2/So400m‑16/512 28.5 41.9 28.5 35.4 46.0
@@ Appendix H Evaluation Method
Consistency check for LLM as a judge
To study potential evaluation bias in our LLM-as-a-judge setup, we additionally evaluate with a different LLM (Qwen3-8B with thinking enabled) as a judge. The results are shown in Table H.1. The results are comparable to those from GPT-4o, with no change in ranking, indicating that cross model validation is consistent.
Table H.1: Evaluating LLM-Judge Evaluation Consistency Qwen3-8B was additionally used as a LLM judge for consistency check.
Model GPT-4o LLM-Judge Qwen3-8B LLM-Judge
o4-mini $55.62_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.509,0.599]}}$ $56.60_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.523,0.613]}}$
gemini-2.5-pro $47.27_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.424,0.519]}}$ $49.54_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.449,0.542]}}$
gpt-4o $26.52_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.225,0.308]}}$ $27.51_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.231,0.322]}}$
Claude-3.7-Sonnet $16.31_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.127,0.199]}}$ $19.91_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}[0.162,0.241]}}$
@@
Using the JSON evaluation log created, we perform repeated sampling with replacement to simulate multiple evaluation scenarios. Specifically, we draw N bootstrap samples of size —S— (equal to the number of total examples), calculate the accuracy for each sample, and estimate the distribution of accuracies. Formally, for each bootstrap sample S with size —S—, the accuracy is computed as:
@@
Details on Human Evaluation Method
We provide details on the set up of human evaluation experiments. The purpose of these evaluations are (1) highlight the gap between VLMs and expert solvers, and (2) provide a detailed error analysis of the failures in VLMs. We recruited several volunteers from our lab - one “expert” (a native English speaker who regularly solves rebus puzzles) and four “non-experts” (one native and three non-native English speakers with little or no prior exposure). Each participant attempted several puzzles (between 25 and 200) that were randomly drawn from the same evaluation set used for the models. The goal with this experiment was not to perform a rigorous comparison between human and model performance, but to give a general idea of how a small subset of humans perform on this task.
@@ Appendix I Prompting
Table I.1: Impact of changing the base prompt on the overall model accuracy (% success).
@@
These models are capable of multi-round chain-of-thought reasoning using special ”thinking” tokens, such as OpenAI o3 and o4-mini OpenAI (2025) models. Gemini-2.5-flash and Gemini-2.5-pro Kavukcuoglu (2025), both from Google, also support this functionality. Notably, Gemini 2.5-flash has both reasoning and non-reasoning variants and we use the reasoning-enabled version in our experiments.
@@
In this section, Figure K.1 shows several qualitative examples from two models (GPT-O3, Phi-4, and Qwen 2.5-VL)and Figure K.2 between models (o4-mini, Claude-3-5-haiku, Phi-4) in our probe dataset.
@@
Figure K.2: This figure shows qualitative examples for both successful (o4-mini) and failed (Claude-3-5-haiku/Last: Phi-4) cases of model prediction.